Two different numbers add to be 54. Three times the smaller number minus the larger number is 14. What is the smaller number?

Answer by math_helper(2461)


x + y = 54 (1) (…and let x be the smaller number)
3x - y = 14 (2)
Add (1) to (2):
4x + 0y = 68 —> x=17 —> (I used eq (1)) —> y=37
—
Ans: 17 is the smaller number
—
Check: 17+37 = 54 (ok)
3(17)-37 = 51-37 = 14 (ok)